OpenStudy (anonymous):

Assume a 125 pound person burns 150 calories after 15 minutes of jogging and 200 calories after 20 minutes of jogging;; Write an equation, in slope intercept form, to represent this data

OpenStudy (anonymous):

Okay, since calories is y, and minutes is x we have two points: (15,150) and (20, 200) Using the slope formula we can find the slope, recall that it is: y2-y1 / x2-x1 200 - 150 / 20 - 15 = 50/5= 10 What does it represent, well slope is rise over run, on this rise you have have calories and run is minutes, so it's calories/minute, or the rate at which calories are burned. For part three we need to make an equation. We have the slope and a point, so it's not to bad. Recall that (y-y1)=m(x-x1) y-200=10(x-20) y-200=10x-200 y=10x Okay, now that we have that we can plug in 35 for x, and find that they will burn 350 calories. :D